# Gravity as a topological gauge theory

### v3

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
In this particular case the fact that Torsion is null implies that the Einstein tensor $G_{kc}$ is symmetric (see [16, p3] for more details). When taking into account that $\eta_{kc}$ is symmetric as well, the anti-symmetricity of $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}$ implies that the energy-momentum tensor $\tau_{ck}$ sports an anti-symmetric part if $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}\neq 0$. This confirms that we can consider a matter Lagrangian that depends on the spin connection $A$ as we explicitly did.
Replacing $e=\dfrac{4\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$; $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ and $y=\dfrac{3}{4}(\dfrac{\pi}{\Lambda_{0}G\gamma}-r)$ in the action, we obtain:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\dfrac{\alpha_{1}}{2}\Bigl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{4}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{(\dfrac{Gr}{\pi}-\dfrac{3}{\Lambda_{0}\gamma})}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{4\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\Bigr{)}$
$\qquad\qquad\overbrace{+\dfrac{\Lambda_{0}r}{12\pi^{2}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}\overbrace{-\dfrac{r}{4\pi^{2}}d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}\biggr{)}.$ (28)
@@
$V(\vartheta)$ $=(\dfrac{2}{\gamma}-\dfrac{2\Lambda_{0}Gr}{3\pi})\alpha_{1}R^{ab}\wedge\beta_{a}\wedge\beta_{b}\overbrace{-\dfrac{\Lambda_{0}\alpha_{1}}{12}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}\overbrace{-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}$ (29)
@@
$\qquad\overbrace{-\dfrac{2r\Lambda_{0}G\alpha_{1}}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{i}\wedge\beta_{j}\bigr{)}}^{Nieh-Yan}\overbrace{+\dfrac{\Lambda_{0}r}{12\pi^{2}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}\overbrace{-\dfrac{r}{4\pi^{2}}d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}.$ (30)
Setting $\lambda=0$ we retrieve the action (16), namely:
@@
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\dfrac{\alpha_{1}}{2}\biggl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{4}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
@@
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{(\dfrac{Gr}{\pi}-\dfrac{3}{\Lambda_{0}\gamma})R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{4\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}.$
@@
This shows that by imposing specific constraints on $\alpha$ and $\lambda$ for $G=SO(4,2)/\{\pm I\}$, one can recover the characteristic polynomial based action for $G=SO(4,1),~{}SO(3,2),~{}\text{or }ISO(3,1)$ with respective constraints:

### v5

@@ 3.2.2 The total action
The secondary source of curvature $-\dfrac{1}{24\alpha_{1}}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\bigl{(}\mathfrak{s}^{ab}_{~{}~{}abr}+(\dfrac{1}{2\pi G\gamma}-\dfrac{\Lambda_{0}r}{3\pi^{2}})T^{m}_{~{}rm}\bigr{)}\varepsilon^{\mu\nu r}_{~{}~{}~{}\,k}$ in (28) is highly dependent on the value of $\Lambda_{0}$ and $r$. One can for example tune $r=\dfrac{3\pi}{2\Lambda_{0}G\gamma}$ such that it is now dominated by the term $-\dfrac{1}{24\alpha_{1}}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\mathfrak{s}^{ab}_{~{}~{}abr}\varepsilon^{\mu\nu r}_{~{}~{}~{}\,k}$.
In case $v=0$ we retrieve the usual equations of motion (11)-(12). We will show in the next section the action $S_{G}+S_{M}$ is asymptotically topological in the limit $\Lambda_{0}\rightarrow 0$.
Replacing $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$; $\ell\ell^{\prime}=-\dfrac{3k}{\Lambda_{0}}$ and $y=\dfrac{3\pi}{4\Lambda_{0}G\gamma}-\dfrac{r}{2}$ in the action (3.2), we obtain:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\bigl{(}rP(\bar{\Omega})+\dfrac{3\pi}{2\Lambda_{0}G}\operatorname{Pf}(\dfrac{F}{2\pi})+(\dfrac{3\pi}{4\Lambda_{0}G\gamma}-\dfrac{r}{2})\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})\bigr{)}$ (31)
$=\int_{\mathscr{M}}\biggl{(}\dfrac{\alpha_{1}}{2}\biggl{(}\overbrace{\underbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}_{Palatini}+(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{\underbrace{-(\dfrac{2Gr}{\pi}+\dfrac{3}{\Lambda_{0}\gamma})}_{\tfrac{2\vartheta}{\alpha_{1}}}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{8\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\Bigr{)}$
$\qquad\qquad\overbrace{-(\dfrac{r}{8\pi^{2}}+\dfrac{3}{16\pi\Lambda_{0}G\gamma})d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}\biggr{)}$
@@
$V(\vartheta)$ $=(\dfrac{1}{\gamma}-\dfrac{2\Lambda_{0}Gr}{3\pi})\alpha_{1}R^{ab}\wedge\beta_{a}\wedge\beta_{b}\overbrace{-\dfrac{\Lambda_{0}\alpha_{1}}{12}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}\overbrace{-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}$
@@
$\qquad\overbrace{-\dfrac{4\Lambda_{0}G\alpha_{1}r}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\overbrace{-(\dfrac{r}{8\pi^{2}}+\dfrac{3}{16\pi\Lambda_{0}G\gamma})d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}.$ (32)
Setting $\lambda=0$, we retrieve the action (15), namely:
@@
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\dfrac{\alpha_{1}}{2}\biggl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
@@
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-(\dfrac{2Gr}{\pi}+\dfrac{3}{\Lambda_{0}\gamma})R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{8\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}.$
@@
In particular, this shows that by imposing specific constraints on the components $\alpha$ and $\lambda$ of the Cartan connection $\varpi$ for $G=SO(4,2)/\{\pm I\}$, one can recover the action for the subgroups $SO(4,1),~{}SO(3,2),~{}\text{or }ISO(3,1)$ already considered in Section 2. Indeed, one has with the respective constraints: